Suppose the probability of a particular stretch of the beach being hit by a hurricane is 1/20.

a) What is the probability of the beach not being hit by a hurricane? _______

b) What are the odds against the beach being hit by a hurricane? _______

c) What is the probability of the beach being hit by a hurricane two years in a row? _______
Answer #1

a) P(Not being hit) = 1 - 1/20 = 19/20

b) Odds against hit = 19 : (20 - 19) = 19 : 1

c) P(Hit two years in a row) = (1/20)*(1/20) = 1/400